OpenStudy (anonymous):

What is the quotient of 4 and 2 over 3Division sign 2 over 3? (WILL FAN AND MEDAL)

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@MrProphecyXII

OpenStudy (anonymous):

can u help me

OpenStudy (haseeb96):

can you write it in equation ?

OpenStudy (anonymous):

\[4\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[\div \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

4 and 2 thirds divided by 2 thirds?

OpenStudy (anonymous):

\[4\frac{ 2 }{ 3 } \div \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

It's 7 I think..

OpenStudy (anonymous):

you can re write the equation as \[\frac{ 14 }{ 3 } \div \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

now you can solve it

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

just simplify it the answer would be 7

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so i gets a medal XD